UT Apr 18 2022 - DEC Night




Night Info

  • Moonrise 20:08
  • Sunset 18:51
  • Nautical Twilight Ends 19:47
  • Astro Twilight Ends 20:18
  • Astro Twilight Starts 04:20
  • Nautical Twilight Starts 04:51
  • Sunrise 05:46
  • Moon illumination 0.98

Staff

  • OSA: J.Williams
  • ISp: C.Bolyard/Mtn
  • Mtn.Manager: J.Prothro
  • ISA/Observer: B.Rothberg, R.T.Gatto, O.Kuhn
  • Special Guests: JC Guerra


Night Plan

  • Starting config LBC-Red
    • LBC Red tasks
    • LBC Extra Focal images in equilibrium -IT 8608, IT 8592 - 1hr
    • Thermal Tracking - TMS Active - need thermal gradient - or create one with adjustment to mirror ventilation -1.5-2hr
    • PHAs recovery - 0.75hr
      • CP for collimaion is at 14:39 +34:28 - target is at 14:03 +33:44 tonight - this is close enough
      • Target Body: (2019 XO2)
    • ********************************************************************************************************************************
      Date__(UT)__HR:MN R.A._____(ICRF)_____DEC dRA*cosD d(DEC)/dt Azi____(a-app)___Elev APmag S-brt Lun_Sky_Brt sky_SNR
      ********************************************************************************************************************************

      2022-Apr-18 02:30 N 14 03 44.91 +33 45 33.2 -71.6213 -19.7317 65.624760 26.898704 23.511 n.a. n.a. n.a.
      2022-Apr-18 03:00 Am 14 03 42.03 +33 45 23.4 -71.8589 -19.8080 68.278121 32.730407 23.511 n.a. 2.241 n.a.
      2022-Apr-18 03:30 m 14 03 39.14 +33 45 13.5 -72.0848 -19.9019 70.760072 38.667513 23.511 n.a. 2.200 n.a.
      2022-Apr-18 04:00 m 14 03 36.25 +33 45 03.6 -72.2950 -20.0122 73.084648 44.692506 23.511 n.a. 2.290 n.a.
      2022-Apr-18 04:30 m 14 03 33.35 +33 44 53.6 -72.4857 -20.1377 75.257944 50.790179 23.512 n.a. 2.396 n.a.
      2022-Apr-18 05:00 m 14 03 30.44 +33 44 43.5 -72.6534 -20.2769 77.273166 56.947137 23.512 n.a. 2.472 n.a.
      2022-Apr-18 05:30 m 14 03 27.52 +33 44 33.4 -72.7952 -20.4279 79.098265 63.151309 23.512 n.a. 2.522 n.a.
      2022-Apr-18 06:00 m 14 03 24.60 +33 44 23.2 -72.9085 -20.5888 80.640363 69.391330 23.512 n.a. 2.554 n.a.
      2022-Apr-18 06:30 m 14 03 21.67 +33 44 12.9 -72.9912 -20.7574 81.615870 75.655217 23.513 n.a. 2.574 n.a.
      2022-Apr-18 07:00 m 14 03 18.74 +33 44 02.5 -73.0416 -20.9314 80.814598 81.924171 23.513 n.a. 2.586 n.a.
      2022-Apr-18 07:30 m 14 03 15.82 +33 43 52.1 -73.0588 -21.1085 61.247783 88.048508 23.513 n.a. 2.591 n.a.
      2022-Apr-18 08:00 t 14 03 12.89 +33 43 41.5 -73.0424 -21.2862 282.823161 85.323573 23.513 n.a. 2.592 n.a.
      2022-Apr-18 08:30 m 14 03 09.96 +33 43 30.9 -72.9924 -21.4620 278.356197 79.076853 23.513 n.a. 2.590 n.a.
      2022-Apr-18 09:00 m 14 03 07.03 +33 43 20.1 -72.9097 -21.6335 278.667358 72.805696 23.514 n.a. 2.584 n.a.
      2022-Apr-18 09:30 m 14 03 04.11 +33 43 09.3 -72.7954 -21.7985 279.960033 66.549621 23.514 n.a. 2.574 n.a.
      2022-Apr-18 10:00 m 14 03 01.20 +33 42 58.4 -72.6515 -21.9546 281.648414 60.322646 23.514 n.a. 2.561 n.a.
      2022-Apr-18 10:30 m 14 02 58.29 +33 42 47.5 -72.4802 -22.0998 283.565545 54.136179 23.514 n.a. 2.543 n.a.
      2022-Apr-18 11:00 m 14 02 55.39 +33 42 36.4 -72.2844 -22.2322 285.653566 48.001789 23.514 n.a. 2.520 n.a.
      2022-Apr-18 11:30 Am 14 02 52.49 +33 42 25.3 -72.0673 -22.3502 287.894303 41.932057 23.515 n.a. 2.487 n.a.
      2022-Apr-18 12:00 Nm 14 02 49.61 +33 42 14.2 -71.8326 -22.4523 290.287749 35.941113 23.515 n.a. n.a. n.a.
      2022-Apr-18 12:30 Cm 14 02 46.74 +33 42 03.0 -71.5840 -22.5374 292.844020 30.045112 23.515 n.a. n.a. n.a.
    • If conditions too poor to continue with LBCs - reconfigure early
  • Reconfigure for Bino-LUCI 00:20
  • LUCI-1 AO/ESM
    • DONE OSCO w/NCPA on - 0.5hr
    • DONE Faint Limit Test WFS (Rmag>16.5) - 1hr - high priority
      • AO Stars FULL DITHERS NEW folder
        • (RA08-09hr Star 10 R=16.6 (17.8) - twlight-00hr)
        • (RA10-11hr Star 5 R=16.6 (17.8) - twlight-01:45hr WD)
        • (RA10-11hr Star 6 R=17.0 (18.2) - twilight-03hr WD)
        • DONE RA12-13hr Star 4 R=16.6 (17.8) - all night WD GD 321
        • RA12-13hr Star 8 R=17.3 (18.5) - 21:15-04hr G-type
        • RA12-13hr Star 9 R=16.7 (17.9) - 20:30-04:30hr WD HS 1338+0807
        • DONE RA14hr-15hr Star 4 R=17.1 (18.3) -20:30-twilight WD SBSS 1510+510
    • Close Separtion Targets test (0.1"-2") - 1-2hr - medium priority (if time)
      • in OT: AO STARS - FULL DITHERS *NEW*-->Binaries
        • Target 2 BD+00 2888 R=10.7 - separation 0.225" - need 256x2048 subframe - visible twilight-02:30hr
          • Target 5 GD 135 R=14.6 - separation 0.3" - no subframe - visible twilight-01:30hr
        • Target 4 SBSS 1148+564 R=14.9 separation 0.6" - no subframe - visible twilight - 04hr
        • Target 6 2MASSJ1128+3136 R=13.5 - 1" separation - no subframe - visible twilight - 03hr
        • Target 7 2MASSJ13061537+2043444 R=8.5 - 1.6" separation -64x2048 subframe - visible twilight-twilight
          • Note: given separation and lack of GS only 1 PA script is available - companion should be visible directly below main star
          • Target 9 2MASSJ10590395+6349283 Bin 2 R=12.7 - 1.57" separation - No subframe backup - visible twilight-03:30hr
        • Target 3 BD+33 2137 R=10.3 - 2.4" separation no subframe
          • Target 10 2MASSJ15553957+4025135 Bin2/3 R=13.3 - 2.3" separation - visible 21:45-twlight
            • Interesting Targets - low priority (if time)
                  • Cat's Eye Nebula - 23hr-twilight
                    • FeII and Hbeta needed (H2 already collected)
                    • Dec +66
                  • DD0 82 - dwarf galaxy w/Bin 1 off-axis star
                    • J,K - twilight-03:30hr
                    • Dec +70
                • ESM Science LUCI-G210 H-band (restframe CaT) spectroscopy of mid-z ULIRG - 1.5hr (if time)
                  • Bin 1 AO Ref star
                  • SL taken in parallel to characterize improvement
              • Reconfigure for MODS-Bino for D-time science by 03:45hr
                • V1674 Her Followup - MODS Bino - M.Wagner - 0.75hr
                • Spectrophotometric star (if clear take all modes) - BD+33

Poor Weather Backup

Eng Tasks
  • Mixed-mods on-sky LUCI+MODS test -0.5hr
    • send active bino preset
  • LUCI Photometric Standards (any seeing, if so bad can't guide, use track mode) - requires clear skies
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
    • Select stars based on visibility, need to get stars at two different airmasses
  • Pointing Model - DX needs improvement after ASM re-install
    • Seeing up to ~2.5"
    • Use LUCIs
      • N375 cameras, N30_FieldStop (aka the "Pointing") masks, K-band, 2.51 sec NDIT=2
      • no flexure compensation (doesn't matter with the pointing mask in)
      • Before taking pointing data, take at least 2 images with dROT=180 for rotator center (I usually do 4)
      • use PtStars in IRAF - current or near current LST - re-run every 2-4 stars with current LST
      • active preset to collimate on-axis, park probe, then take image - 1st image
      • turn PCS logging ON (it is possible this is always on now...check)
      • send track mode presets to stars after that Note: if you use off-axis stars and let it collimate at each point, you can also make a collimation model
      • Turn PCS logging off, note filename
Science
  • Rothberg - Local Mergers
    • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V43.xml
    • Not in order of priority
      • NGC 4004 - 21hr-twilight Dec +27
      • NGC 3597 - 23hr-02hr Dec -23
      • UGC 9829 - 01:30hr-twilight Dec -01
      • UGC 10675 - 02hr-twilight Dec +31

Night Log

Closed Dome

  • UT 00:17 - Init All LUCIs, bringing up GUIs, test dark
    • L1 0001-0002 , L2 0001-0002 - 2.51sec dark
  • UT 00:27 - MODS awake - simSnap OK
  • UT 00:28 - check N30Fieldstop - putting N30Fieldstop in and switching to N30 camera
  • LUCI-1 Camera Wheel Filter Error again - bringing up eningeering panels
    • switch 11 and 12 are off in Hiramo, unknown camera positoin on MCU is 198
    • initializing camera wheel
    • at N375, -602 position and switch 12 is on
    • moving to N30, Hiramo says switch 11 and 12 are on, at position +27 in MCU for camera
    • updated IT 8183
  • N30 Fieldstop check L1 0003-0006 - OK
  • O.Kuhn bringing up LBC-Red

On-Sky

  • Copied reference from 20220408 to /tmp - ran tms_loop_right.pl, but POSE calculations failing on both sides. Channels 1,2,5,12,13 looked good on "Alignment" mode - others < about 2%
  • Tried editing tms_loop_right.pl to change --mcp none to --mcp dx. That didn't help. There is an error, but probably unrelated.
    Traceback (most recent call last):
      File "/lbt/lbto/supportscripts/MetrologySupport/lbc_tms.py", line 739, in <module>
        sx_x=float(result[0].strip("[(,)]"))
    ValueError: could not convert string to float: '*'
    RIGHT ONLY.... Sleeping for 5 seconds - ^C to exit
    /lbt/lbto/supportscripts/MetrologySupport/lbc_tms.py script version of 12-Apr-2022
  • Finally I realized - there are probably only 5 useful channels, not the minimum of 6 needed to compute a pose - is that correct? What happened? The DX reference from 20220408 used channels 1,4,5,6,10,12,13.
  • I turned the laser off and back on. There was some oddity in the start up. This time, when turning it on, there was a popup "to call a function both the library name and function are needed." I think I saw this before. Now, however, when I checked the alignment, I see many more channels that are above 5%.
  • 02:47 Run active loop again (after restoring mcp to none). Now TMS is making and sending corrections.
Thermal extra-focal pupil test
  • 02:48 Slewing to field at 10:30 20deg. Waiting for TMS update.
  • PMC: glass = 6.63, ambient 8.193
  • dofpia. I messed up my copointing script and was pointing at a random field with only 3 stars.
  • Fixed it and ran dofpia again. This time there are 8 stars.
  • Injected z4/z11
  • UT 03:08 ptautoadjust on red - pointing correction about 10"
  • Ran tms_loop_right_x.pl:
    • tms_loop_right_x.pl is a copy I made of tms_loop_right.pl where I added the option "-x" to the lbc_tms.py command line (after "-s 1")
  • UT 03:20 start taking pupil data. They are already comatic! (There is a good reason - I forgot to set the reference).
  • Run dofpia, /backout and also ./remsph.py -s 1 to remove z4/z11 that had been injected by FPIA and manually.
  • Run dofpia again. Set reference. oops - I forgot to inject spherical first. Add z4/z11 via "addsph.py -s 1". Set reference again. Now start the tms_loop_right_x.pl and run the OB.
  • 03:30 Starting to take pupil data. Glass=7.1 Ambient = 8.2
  • There are at least 2 extra-focal star images that have an odd shape or a dark curve running through - too wide for a detector artifact. These are at approx (1715, 1700) and (1885, 1500) on lbcr.20220418.034914.fits. These two stars are too bright for FPIA anyway. I checked a few other pupil images, but I don't see irregularities on these.
  • 04:28 Glass=7.686 Ambient 9.000
  • Ran FPIA offline on the images: z4 is consistently about 1000 nm, from the start. That is odd [so maybe I should not have injected z4/z11 before setting the reference - these values went into the Zernike global offset registers]. But I see coma growing - z7 and z8. The Z11 and Z22 estimates are also showing some trends.
  • LBTplot --> dxTMS is not working. There is an error about not being able to access /tmp/*_left_tmsoffset.dat.
[telescope@robs ~]$ LBTplot
ls: cannot access '/tmp/*_left_tmsoffset.dat': No such file or directory
LBTplot: Process timeout from  tms_right  at  04:49:06  for  160  sec
  • 05:00 DIMM is reporting 0.93 (zen-corr). The target is at elev 73 deg.
  • 05:24 Glass = 8.113 and Ambient = 8.685
  • 05:25 Stopping OB and stopping TMS active
  • remsph.py -s 1 to remove the z4/z11 offsets
  • dofpia, /redonly (noticed that FPIAs z4/z11 offsets go indto a different register).
    • 1438 -62 151 952 -743 -280 -71 are the first offsets.
    • On the converged pupil images, there is no sign of the bright ring.

  • set the reference
  • inject z4/z11
  • set the reference again (in case there is a difference)
  • run tms_loop_rihgt_passive.pl
  • 053500 delta_Z22 = +150
  • 053744 delta_z22 = -150
  • 053946 delta z11 = +1500 delta z22 = 0
  • 054147 delta z11 = -1500 delta z22
  • 054344 delta z22 = +300 delta z11 = 0 - see rings in inner & outer edge - subtle
  • 054534 delta z22 = -300 delta z11 = 0 - see bright ring around middle developing
  • 054724 delta z22 = -500
  • 054852 delta z22 = +500 - right PSF bad new influence forces... looks like move may have failed?
  • 055128 delta z22 = +400 (no error now on PSF)
  • With passive loop running still, taking out the forces: z22 by hand and z4/z11 by remsph.py -s 1
  • dofpia, /redonly to see if there was any drift.
    • There was some drift in z4: -466 37 83 -12 2 12 -21 are z4-z11,z22
  • setting reference
  • starting tms active

2019 XO2 observation
  • 05:58 Slew to pointing field for 2019 XO2 (new one created with more stars and closer to target).
  • 06:02 Playing the script to check pointing after having seen one TMS correction come through. Is IQ ok?
    • Ran ptautoadjust to correct pointing. A 28" offset was needed to IE: dCA=0.0, dIE=28.1
    • IQ is good --- the pointing star looks round and has FHWM 2.34 pixels!
  • 06:09 Starting the non-sidereal observation.
    • I forgot to turn off guiding! Pausing the OB after the current exposure and I'll turn off guiding through the user interface. Guide star has FWHM 0.8". Image is good.
      • But, chaos... I turned off guiding and resumed from the pause, but forgot to set the NS hijack and we slewed a long, long way.... (RA, DEC = 11, 06) but the target is at 14, 30.
    • 06:19 Taking the next 2 exposures of the script, NS (clicked set) and not guiding (turned off guiding in UI). This will be a good test of TMS - (for the one untrailed object)!
  • Taking a sidereal exposure on the CP field to check IQ. FWHM is about 2.3-2.6 pix and the images look nice. I see a bit of aberration but this is LBCR chip2 for which there is a PSF gradient.
Mixed Mode
  • During this test, I kept TMS running in active mode on DX.
  • Josh is sending a pointing and collimation preset for LUCI1 in pseudo-mono mode.
  • By 06:58 we were pointed and collimated.
  • Since we had already collimated with the LBCR, we skipped this step and just sent the synch presets from LUCI1+LBCR.
  • First LBCR exposure is 065837.
    • We did not see any problems.
    • We took 3 LBCR exposures and several LUCI dithered exposures in all.
lbcr.20220418.065837.fits[2] 177-000187 14:02:37.086 +34:23:52.37 r-SLOAN 120.241 -179.59461 0 0
lbcr.20220418.070124.fits[2] 177-000187 14:02:37.248 +34:23:56.37 r-SLOAN 120.239 -179.83720 2 4
lbcr.20220418.070422.fits[2] 177-000187 14:02:36.925 +34:23:48.37 r-SLOAN 120.240 -179.75673 -2 -4
2019 XO2 observation
  • IQ looked ok on last LBCR image on copointing field. The FWHM had gone up to 3.23 FWHM.
  • When Josh switched from LUCI+LBCR to just LBCR, we hit limits on DX. The pointing offsets were about 1.2mm in X and this, plus the the lookup table sent the mirror to the limits. When I sent the LBCR preset, the pointing offsets all became 0 and the mirror went out of the limits. All during this time, I was running TMS active.
  • I realized that I ran the script I made up for testing Mixed-Mode, but with the NS hijack. So this took 2 120-sec exposures at PA=180 in r-SLOAN.
  • 07:20 Now running the original 3 x 180-sec script.
  • The set of images below are all non-sidereally tracked on 2019 XO2, including the two with object=177-000187, PA=180 and exptime=120s.
lbcr.20220418.060808.fits[2] 2019 XO2 14:03:23.824 +33:44:20.49 r-SLOAN( 180.239 90.06411 0 0
lbcr.20220418.061925.fits[2] 2019 XO2 14:03:22.056 +33:43:56.62 r-SLOAN 180.239 90.04719 -10 -20
lbcr.20220418.062311.fits[2] 2019 XO2 14:03:23.025 +33:44:35.33 r-SLOAN 180.240 90.05126 10 20
lbcr.20220418.071115.fits[2] 177-000187 14:03:17.930 +33:43:56.69 r-SLOAN 120.239 -179.82795 4 -2
lbcr.20220418.071357.fits[2] 177-000187 14:03:17.134 +33:43:59.74 r-SLOAN 120.240 -179.82726 -4 2
lbcr.20220418.071909.fits[2] 2019 XO2 14:03:17.001 +33:43:56.31 r-SLOAN 180.239 90.93804 0 0
lbcr.20220418.072250.fits[2] 2019 XO2 14:03:15.866 +33:43:34.64 r-SLOAN 180.240 90.22009 -10 -20
lbcr.20220418.072638.fits[2] 2019 XO2 14:03:16.838 +33:44:13.35 r-SLOAN 180.239 90.35177 10 20

  • 07:35 Stopping TMS and turning off the laser.

AO Testing

  • UT 07:49 - pointing check
OSCO
  • Target is BD+08 2833 Rmag=7.48, WFSMag=8.7
  • UT 07:58 - ASM is in failure - SX RIP'd
  • doing AdSec start and stop
  • J.C.Guerra noted that TSS was activated - but looks like one data point of crossing the TSS threshold in LBTplot - otherwise winds are very low tonight
  • UT 08:03 - setting shell
  • UT 08:04 - resending preset
  • NCPA enabled during first exposure
  • First pos. telemetry w/ NCPA:

    [telescope@obs1 ~]$ ./get_ao_telemetry

    => Command line arguements processed
    Set 0 Tracknum 20220418_081016 LuciName unknown
    Set 1 Tracknum 20220418_081029 LuciName unknown
    Set 2 Tracknum 20220418_081042 LuciName unknown
    -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220418_081016.log
  • UT 08:14 - SX ASM failure - no TSS trigger - OCAM went not live
  • 2nd pos. telemetry w/ NCPA:

    [telescope@obs1 ~]$ ./get_ao_telemetry

    => Command line arguements processed
    Set 0 Tracknum 20220418_081250 LuciName unknown
    Set 1 Tracknum 20220418_081303 LuciName unknown
    Set 2 Tracknum 20220418_081315 LuciName unknown
    -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220418_081250.log
  • JC.Guerra reports TSS is activated - it shouldn't have been
  • UT 08:22 - shell is flat again
    • L1 0010-0027
    • #0028 affected by failure
  • JC Guerra reports fast diagnostic disconnected
  • UT 08:23 - once SX ASM is back we will run OSCO again - JC.Guerra is worred about stablitiy - so running a Bin 1 star again
    • Guider shows 0.55"
  • Is the seeing too good for SOUL?
  • NCPA on
  • 1st pos. telemetry w/ NCPA:

    [telescope@obs1 ~]$ ./get_ao_telemetry

    => Command line arguements processed
    Set 0 Tracknum 20220418_083035 LuciName unknown
    Set 1 Tracknum 20220418_083047 LuciName unknown
    Set 2 Tracknum 20220418_083059 LuciName unknown
    -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220418_083035.log
  • 2nd pos. telemetry w/ NCPA:

    [telescope@obs1 ~]$ ./get_ao_telemetry

    => Command line arguements processed
    Set 0 Tracknum 20220418_083325 LuciName unknown
    Set 1 Tracknum 20220418_083337 LuciName unknown
    Set 2 Tracknum 20220418_083349 LuciName unknown
    -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220418_083325.log
  • UT 08:35
    • L1 0029-0048
Faint Star Test
GD 321 - White Dwarf Rmag=16.6
  • UT 08:37 - preset sent
  • Rmag 16.6, WFSmag = 17.7
  • Hmag=16.1
  • 74 deg from Moon
  • Elevation = 53deg
  • PA240, Instrument Rotator=132.6
  • On preset Guider showed 0.83", now (UT 08:46) DIMM shows 1.3"
  • UT 08:47 - loops closed - LUCI starting
  • N30 Fieldstop really shifted - but not critical to this test
  • Frame #0049 - FWHM= 0.25"
  • image quality continues to be poor - loop running 100Hz
  • UT 08:52 - loop opened - safeskip
  • UT 08:54 - DIMM 1.1", guider 1.1" - 2photons/subapp
    • L1 0049-0052
  • UT 08:54 - resending preset - failed - cannot find start on CCD47
  • 2022.04.18 08:55:01 s | error high | luci.luc | LBT | TelescopeServer | RMITelescopeImpl.java#iifCommandMessage(1439) | IIF PresetTelescope command for left side(s) failed! 0- PresetTelescope result status: Error 1- Stop{FLAO}[1013] 2- failed to find a star in image
  • Could be pointing - Guidestar was on the edge
  • UT 08:56 - pointing check
  • UT 08:58 - sending preset again
    • guider shows 0.89"
    • star is visible in CCD47
    • not applying GOpt
  • UT 09:02 - loops closed
  • WFSmag = 17.7, Rmag =16.6
  • UT 09:06 - taking LUCI data
    • Frame #0054 - FWHM=0".25
  • UT 09:09 DIMM ~1"
    • Frame#0057 - FWHM=0".25
  • UT 09:16 - appeared back on CCD47 - was slightly off but took a few seconds to correct the centering
    • Would be useful if the centering occurred before handing back off to LUCI
  • UT 09:20 - Seeing 1.4" on DIMM
    • L1 0053-0068
SBSS 1510+510 White Dwarf Rmag=17.1
  • UT 09:29 - sending preset
  • Hmag 17.0
  • Rmag=17.1, WFSmag =17.96
  • PA 120, instrument rotator = -23.7
  • UT 09:31 - re-rotator out of range
    • using MCGUI to move to 200 deg on rotator then resending the preset
  • UT 09:33 - sending preset, instrument rotator = 338
  • Elevation = 70
  • guider shows 0.6"
  • Can see the star in CCD47, 1 photon per subapp
  • UT 09:41 - running GOpt
  • UT 09:42 - failed on apply optical gain, clicking done
  • UT 09:42 - starting LUCI exposures
    • Frame#0069 - FWHM =0"36
    • Frame#0071 - FWHM=0".35
    • N30Fieldstop shifted again
    • Frame#0078 - FWHM = 0".36
    • Frame#0080 - FWHM = 0".25 - strehl ~1%
    • L1 0069-0084
  • UT 10:05 - done
  • This is the star on CCD47
  • Screen Shot 2022-04-18 at 3.05.33 AM.png
UT 10:09 - aligning N30Fieldstop L1 0085-0087

Cat's Eye Nebula Part Deux

  • UT 10:12 - sending preset
  • will take H2 image then skip to FeII
  • Rmag=11.38, WFSMag =13.2
  • Instrument Rotator=159
  • UT 10:22- starting data
  • Starting with FeII filter
  • N30 Fieldstop shifted again - had just fixed it
    • L1 0089-0090
    • L1 0088- short exposure - offset fieldsop
  • updated FeII and H2 to 20sec x 3 NDIT
  • UT 10:29 -restarting
    • FWHM=0".061
  • LUCI READOUT GUI is in a funky state
    • Frame#0092 - saved a single image - looks like it jumped
    • Freme#0091 - saved as single frame but looks OK
  • UT 10:47 - DIMM 0.96"
  • NCPA off for all of FeII
    • L1 0093-0105 - FeII NO NCPA
  • UT 10:50 - JC Guerra turning on NCPA
    • L1 0106-0117 -Pbeta NCPA ON
    • PaBeta FWHM=0".058
  • Cat's Eye Nebular in CCD47
  • Screen Shot 2022-04-18 at 3.30.58 AM.png
UT 11:10 - authorizing Binocular LUCI - running ESM

UT 11:11 - running pointing and collimationccd

UT 11:12 - OBS5 froze on me

UT 11:15 - sending pointing

Ring Nebula M57

  • UT 11:22 - sending preset to M57
  • guiders show 0.8"-0.9" both sides
  • Rmag=12.2, WFSmag=14.1 Instrument Rotator=94.8
  • UT 11:33 - starting exposures
  • L1 FWHMs 0.32" AO Ref star is 0.4"
  • UT 11:47 - guidestar vanished on SX but AO is running - its there on right -50, +15 offset
  • Frame#0130 - AO Ref star = 0.3" same star on L2 0.56"
    • L1 0118-0133
    • L2 0005-0020
UT 12:00 - closing up - will take cals

UT 12:01 - putting MODS to sleep, turning off LBC-Red

Closed Dome Cals

  • N375 Bino H2
    • L1 0133-0142
    • L2 0020-0029
UT 12:21 - putting LUCI-2 in safe config
  • N30 H-band, FeII, and PaBeta
  • N30 Fieldstop shifted yet again - first set of FeII don't use - resetting it
    • skip L1 0143-0156
    • L1 0157-0166 FeII
    • L1 0167-0176 PaBeta
    • L1 0177-0186 H-band
UT 12:49 - putting LUCI-1 in safe config

DONE

Summary of LBC work

Only LBCR was available, but we nevertheless ran two LBC/TMS or image quality engineering tests, a D time PHA observation with TMS and made a quick confirmation that the LUCI1+LBCR mixed mode observing works, giving no trailed LBCR images.

Thermal Test: We started the night with the mirror over a degree below the ambient temperature at the cell, and we collimated LBCR and collected almost 2 hours of extra-focal pupil images with z4/z11 injected and TMS running actively. I analyzed a few offline and found that z7 and z8 both increased and there were some trends the z11 and z22 coefficients although they remained fairly small. I also noted that FPIA detected ~1000 nm from the start, and maintained this. I am not sure how this crept in, and after the test ended, I removed z4/z11, stopped the active mode and ran dofpia as usual, and it also detected z4 ~ 1000 nm on the first extra-focal pupil image.

Z11/Z22 grid: After having recollimated, I again injected z4/z11 and obtained a set of pupil images with Z22 = {+150, -150, +300, -300, +500(error), -500, +400} and Z11 = {+1500, -1500}. The results are below. The effects of Z22 of a magnitued ~150-300 are subtle and the pupil image with the Z22 = -500nm offset looks most similar to those having the bright ring that occurs with the Z22 runaway. When this bright ring appears, a large Z22 correction, +500 or more, may be needed. Z11/Z22 pupil grid

Mixed-mode LUCI1+LBCR: We sent synchronous presets from both sides with the LBC SYNCOFFSET tag added to command asynchronous LBCs offsets. The LUCI script dithered by with a 20" throw and LBC by ~2-3". We took 3 LBCR images and a few LUCI images with no problems.

2019 XO2: This went pretty well - we collected two rounds of the 3-dither script plus an extra 2 120-sec images at PA=180 by accident.

-- BarryRothberg - 17 Apr 2022
I Attachment Action Size Date Who Comment
20220418_z11z22grid.pngpng 20220418_z11z22grid.png manage 1 MB 18 Apr 2022 - 08:44 OlgaKuhn LBCR extra-focal pupils with Z11 and Z22 offsets
Screen Shot 2022-04-18 at 3.05.33 AM.pngpng Screen Shot 2022-04-18 at 3.05.33 AM.png manage 56 K 18 Apr 2022 - 10:09 BarryRothberg  
Screen Shot 2022-04-18 at 3.30.58 AM.pngpng Screen Shot 2022-04-18 at 3.30.58 AM.png manage 36 K 18 Apr 2022 - 11:16 BarryRothberg  
Topic revision: r10 - 18 Apr 2022, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback